Suppose a researcher goes to a small college of 200 faculty, 12 of which have blood type O-negative. He obtains a simple random sample of n = 20 of the faculty. Let the random variable X represent the number of faculty in the sample of size that have blood type O-negative.

(a) What is the probability that 3 of the faculty have blood type O-negative? (Round your answer to the nearest thousandth).    

(b)  What is the probability that a none (x = 0) of the faculty has blood type O-negative? (Round your answer to the nearest thousandth).


1
Expert's answer
2021-06-09T07:39:11-0400

"n=20 \\

p=\frac{12}{200}=0.06 \\

q=1-0.06=0.94 \\

(a) \; P(X=3) = \frac{20!}{3!(20-3)!}(0.06)^3(0.94)^{20-3} \\

= \frac{6840}{6} \times 0.000216 \times 0.349279 \\

=0.086 \\

(b) \; P(X=0) = \frac{20!}{0!(20-0!}(0.06)^0(0.94)^{20-0} \\

= 1 \times 1 \times 0.94^{20} \\

= 0.290"